OpenStudy (rosales0228):

From the sum of 5 – 9i and -4 + 7i, subtract 14 – 3i. -13 + i -13 – 5i 13 – i 13 + 5i 13 – 5i

OpenStudy (anonymous):

\[5−9i+(−4)+7i\]

OpenStudy (anonymous):

\[-90 + 7 i \]

OpenStudy (rosales0228):

that's not the answer

OpenStudy (owlcoffee):

One very important thing to know about the complex numbers is that they are represented as vectors, by that I mean they have some special things when it comes to operating them. Say we have two complex numbers: \[u=a+bi\] \[v=a'+b'i\] Now, I want to know the sum of the numbers "u" and "v", it would look like this: \[u+v=(a+bi)+(a'+b'i)\] The way to do this is getting rid of the parenthesis and agroup the real part (the number without "i") and the imaginary part (the number with a "i"). So the sum would look like this: \[u+v=(a+a')+(b+b')i\]

OpenStudy (anonymous):

I'm getting -13-5i ---> 5-9i-4+7i-14-3i = -13-5i